There are two possible methods for recovering copper from a solution of copper (II) sulfate, one uses precipitation and the other uses redox.

Copper(II) ions can be precipitated as copper(II) carbonate.
CuSO4(aq) + Na2CO3(aq) à ƒ ƒ ƒ ¢ † ’ CuCO3(s) + Na2SO4(aq)
1.) How many mL of 0.35 M Na2CO3 would be needed to precipitate all of the copper ions in 26 mL of 0.43 M CuSO4?
2.) How many grams of CuCO3(s) would be formed from the copper ions in 55.4 mL of 1.71 M CuSO4?
1
Expert's answer
2021-10-20T01:59:02-0400

Molar mass of CuSO4= 159.609g/mol

159.609× 0.43= 68.632

26/68.632= 0.379mol

Molar mass of Na2CO3= 105.9888 g/mol

0.379× 105.9888= 40.152g

40.152× 0.35= 14.05ml